OpenStudy (anonymous):

Differentiate. y = (r^2 − 9r)e^r

OpenStudy (valpey):

Product rule: \(\frac{d(uv)}{dr} = u*d{v}+v*du\) \(u=(r^2-9r)\); \(du=(2r-9)\) \(v=e^r\); \(dv=e^r\)